Last year, X Company sold 62,500 units of its only product for $18.00 each. Total costs were as follows:

Cost of goods sold
   Variable $443,125
   Fixed 130,000
Selling and administrative
   Variable $90,000
   Fixed 65,625


At the end of the year, a company offered to buy 4,680 units of the product but only for $11.00 each. X Company had the capacity to produce the additional units, and even though there would have been no additional selling and administrative costs, it rejected the offer.

5. If X Company had accepted the special order, firm profits would have increased by?

6. X Company's production manager thought that the special order units might have required additional direct material and direct labor costs per unit of $0.78 and $0.21, respectively, and the rental of special equipment for $1,500. The combined effect of these changes would have been to reduce the special order profit by?

7. X Company's marketing manager felt that in order for regular sales to continue at 62,500 units this year, the company would have had to reduce the regular selling price this year to $17.43. The result of this price reduction would have been to reduce X Company's profits this year by?
Answer #1

Solution 5:
Variable cost of goods sold per unit = $443125 / 62500 = $7.09 per unit
If X company accept the special order then profit will increase by = ($11 - $7.09) * 4680 = $18,299

Solution 6:
Reduction in special order profit by = ($0.78 + $0.21) * 4680 + $1,500 = $6,133

Solution 7:
Price reduction would reduce X company profit by = 62500*($18 - $17.43) = $35,625